In the US, lengths are often measured in inches, feet, yards and miles. Let's do some conversions. The definition of the inch is: 1 inch = 25.4 mm, exactly. A foot is 12 inches and a mile is 5280 ft, exactly. A centimetre is exactly 0.01 m or 10 mm.

Sammy is 5 feet and 5.3 inches tall. Mark all the true statements below. Use the appropriate number of sig figs.):

A. Sammy is 1.66 m tall.
B. Sammy is less than 0.5 m tall.
C. Sammy is between 1600 and 1700 mm tall.
D. Sammy is more than 168 cm tall.
E. Sammy is 55.3 inches tall
F. Sammy is more than 60 inches tall
G. Sammy is 5.44 feet tall.
H. Sammy is 5.4 feet tall.
I. It would take about 90 000 Sammies, lying head to toe to span the entire Victoria's 90 kile beach.
J. It would take about 100 000 Sammies, lying head to toe to span the entire Victoria's 90 kile beach.
1
Expert's answer
2020-02-26T10:12:28-0500
"x=5 ft \ 5.3\ in=5(12)+5.3=65.3\ in"

"x=\frac{(65.3)}{12}=5.44\ ft"

"x=65.3(25.4)=1660\ mm=166\ cm=1.66\ m"


"N=\frac{90\cdot 5280}{5.44}=90000"

Answer: A, C , F, G, I.